OpenStudy (anonymous):

I need help.. derivative y=4sin^(-1)(3x) I have so far.. y'=d/dx[4arcsin(3x)] y'=4*d/dx(arcsin(3x))*d/dx(3x)

OpenStudy (anonymous):

y = Asin(ωx - φ) Where: y = vertical position A = amplitude ω = angular fequency x = horizontal position φ = phase Period = T = 1/f = 2π/ω

OpenStudy (anonymous):

^wth?

hero (hero):

lol

zepdrix (zepdrix):

Do you know the derivative of arcsine? Are you allowed to use the shortcut, or do you need to show steps on how you got through it? :D

OpenStudy (anonymous):

I don't have to show "exact" steps.. the derivative of arcsine is 1/sqrt(1-x^2)

zepdrix (zepdrix):

Oh ok :) Then it looks like you have it pretty much figured out, only one thing to be careful about though...

OpenStudy (anonymous):

but its arcsine of (3x) how do i do that?

zepdrix (zepdrix):